OpenStudy (anonymous):

Multiply. Try to recognize the type of product before multiplying? (t^2-0.2)(t^2+0.2)

OpenStudy (anonymous):

it's a product of squares

OpenStudy (anonymous):

did not say

OpenStudy (anonymous):

should just be t^4 - .04

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i wasn't asking...i was telling you it's a product of squares

OpenStudy (anonymous):

yes...that's the answer

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanks for the medal?

OpenStudy (anonymous):

(t^2-0.2)(t^2+0.2)= (.t^4-0.04 )

OpenStudy (anonymous):

can i put it like this

OpenStudy (anonymous):

no just put the answer...the way you have it written it's like equaling 1

OpenStudy (anonymous):

(t^2-0.2)(t^2+0.2)= (t^4-0.04 ) so you are correct

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thx for the medal math...now if only the person asking the questions appreciated the effort enough to give them :(

OpenStudy (anonymous):

well thanks for help yall

OpenStudy (anonymous):

your welcome...good luck

OpenStudy (anonymous):

thanks